OpenStudy (anonymous):

The following sum 1/(1+6/n)(6/n)+1/(1+12/n)(6/n)+1/(1+18/n)(6/n)+...+1/(1+6n/n)(6/n) is a right Riemann sum for a certain definite integral \[\int\limits_{1}^{b}\] ∫f(x)dx using a partition of the interval [1,b] into n subintervals of equal length. the integrand must be the function f(x) =

OpenStudy (anonymous):

I found b to be 7. but I can't figure out integrand

OpenStudy (snowsurf):

I figure out f(x) to be \[f(x)=\frac{ 1 }{ (1+\frac{ 6x}{ n }) }\] If the limit starts from 1 to b if the limit was from 0 to b then it would be \[f(x)=\frac{ 1 }{ (1+\frac{ 6x+6 }{ n } )}\]